Alternative Comparison-Different Lives 5.17 Dexcon Technologies, Inc. is evaluating two alter- natives to produce its new pla   

can you solve it by using Excel
Answer #1

Analysis will be carried out for 4 yrs (LCM of 2 & 4)

Using NPV function in Excel

DDM LS
Year Investment Maint Cost Salvage value Net cash Flow Investment Maint Cost Salvage value Net cash Flow
0 -164000 -164000 -370000 -370000
1 -55000 -55000 -21000 -21000
2 -164000 -55000 0 -219000 -21000 -21000
3 -55000 -55000 -21000 -21000
4 -55000 0 -55000 -21000 30000 9000
NPW -420269.3 -409895.8

As the net present cost of LS is lower, it should be selected

Showing formula in excel

DDM LS
Year Investment Maint Cost Salvage value Net cash Flow Investment Maint Cost Salvage value Net cash Flow
0 -164000 =P13+Q13+R13 -370000 =T13+U13+V13
1 -55000 =P14+Q14+R14 -21000 =T14+U14+V14
2 -164000 -55000 0 =P15+Q15+R15 -21000 =T15+U15+V15
3 -55000 =P16+Q16+R16 -21000 =T16+U16+V16
4 -55000 0 =P17+Q17+R17 -21000 30000 =T17+U17+V17
NPW =NPV(20%,S14:S17)+S13 =NPV(20%,W14:W17)+W13